A company that operates for 18 hours a day has introduced a new daily wage system that

incorporates variable hourly rate to encourage longer shifts. The company pays $10 for first

hour or any part of the first hour. Afterwards, company pays additional 2% per hour for each

hour or part of each additional hour. Write a piece-wise function that shows the salary as a

function of hours. Also sketch a graph of this function.


1
Expert's answer
2021-09-27T04:03:13-0400

"2. \text{The piecewise function can be given by }\\
f(x) = 10 \text{ if $0 < x \leq 1 $}\\
f(x) = \frac{2}{100}.10(x-1)+10 \text{ if 1$\leq x \leq 18$}\\
\implies f(x) = \begin{cases} 10 \text{ if 0 < x $\leq$ 1 }\\
\frac{2x+98}{10} \text{ if $1< x \leq 18$}
 \end{cases}\\"